OpenStudy (christian_10_):

Choose all answers that are correct. 2. (1 pt) Which expressions show how 8 • 54 can be rewritten using the distributive property? Choose all answers that are correct. A. 8 • 50 + 4 B. 8 • 50 + 8 • 4 C. 8 • 60 – 8 • 6 D. 8 • 60 + 8 • 4

OpenStudy (anonymous):

b and d

OpenStudy (anonymous):

b and c*

OpenStudy (christian_10_):

Thanks!!!!!!!!

OpenStudy (anonymous):

b and c

OpenStudy (anonymous):

welcome

OpenStudy (mathstudent55):

Original expression: 8 • 54 8 • 54 = 432 The original expression evaluates to 432. A. 8 • 50 + 4 = 404 B. 8 • 50 + 8 • 4 = 400 + 32 = 432 C. 8 • 60 – 8 • 6 = 480 - 48 = 432 D. 8 • 60 + 8 • 4 = 480 + 32 = 512 Since choices A and B equal numbers different from 432, they can be eliminated. Choice B and C evaluate to the correct number, 432. By looking at choices B and C, we see that they both show correct usages of the distributive property, so they are the answers.